OpenStudy (anonymous):

Find the value of y so that the line passing through (2, 6) and (1, y) has a slope of 5. A. B. 9 C. -1 D. 1

OpenStudy (anonymous):

slope = (y2 - y1)/(x2-x1) 5 = (y-6)/(1-2) 5 = y-6/-1 -5 = y - 6 -5 +6 = y

OpenStudy (saifoo.khan):